OpenStudy (anonymous):

complete the square x^(2)+(3)/(4)x+__=(x+__)^(2)

OpenStudy (mertsj):

1/2 of 3/4 = 3/8 and 3/8 squared is 9/64 so add 9/64 and the factors are (x+3/8)^T2

OpenStudy (mertsj):

I meant (x+3/8)^2

OpenStudy (anonymous):

but wat about the other number ???

OpenStudy (mertsj):

Did I say to add 9/64?

OpenStudy (mertsj):

x^2 +3/4x +9/64 = (x+3/8)^2

OpenStudy (anonymous):

yup that was it thnks